How would you rewrite the formula A=1/2(b1+b2)h to find b2?
Answer by Edwin McCravy(20056)

How would you rewrite the formula A=1/2(b1+b2)h to find b2?

               1
          A = ---(b1 + b2)
               2

Multiply both sides of the equation by 2 to clear of fractions:

                 1 
        2·A = 2·---(b1 + b2)h
                 2

Cancel the 2's on the right
     
              1  1 
        2·A = 2·---(b1 + b2)h
                 2
                  1

         2A = (b1 + b2)h

Since mosdt people like to distribute
with the outer multiplier on the left
so we rewrite the right side:

         2A = h(b1 + b2)

         2A = hb1 + hb2

Isolate the term in b2 by subtracting 
hb1 from both sides.

   2A - hb1 = hb2

Divide both sides by h

  2A - hb1
 ---------- = b2
     h 

Edwin